OpenStudy (anonymous):

Evaluate 4b + 2c if b = 3.5 and c = 9

OpenStudy (anonymous):

Substitute the values: i.e.\[4b+2c = 4(3.5)+2(9) = 14 + 18 = 32\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait I don't understand...

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

can you like go over it please...

OpenStudy (anonymous):

The expression is\[4b+2c\] Its given, b=3.5 and c = 9. So the find the value of (4b+2c) just put the values of b and c. \[4 \times 3.5 + 2 \times 9\] Got it?

OpenStudy (anonymous):

oh......

OpenStudy (anonymous):

okay I got it thanks!!!

OpenStudy (anonymous):

welcome.